OpenStudy (anonymous):

How do you solve something like this: 16!/4!12! (The 4!12! are underneath the 16 though. So it looks like a fraction.)

OpenStudy (anonymous):

You can start expanding it:\[{16! \over {4!12!}} = {{16 \times 15 \times 14 \times 13 \times 12!} \over {4! 12!}}\]Then cross out the 12! and... you can solve the rest ;)

OpenStudy (anonymous):